Write an equation of the line that passes through the given point and has the given slope m. (5,1);m=2
Found 2 solutions by stanbon, chitra:
Answer by stanbon(75887)

Write an equation of the line that passes through the given point and has the given slope m. (5,1);m=2
------------------------------
Line Form: y = mx+b
You are given m=2,x=5,y=1; solve for b"
1 = 2*5 + b
b = -9
--------------
EQUATION:
y = 2x -9
================
Cheers,
Stan H.

Answer by chitra(359)

The equation of the line passing through the point (5,1) and with slope = 2 is given by:

(y - y1) = m(x - x1)

(y - 1) = 2(x - 5)

y - 1 = 2x - 10

y - 2x -1 + 10 = 0

y - 2x + 9 = 0

Thus the solution.

regs